I need to know the positive factors for the following:: 32,37,53,108, and 157. i have 1 and the number but i don't know how to get the other ones! Please help me!
Answer by BGluvsmath(64)

32
Solved by pluggable solver: Factor any number
32 is NOT a prime number: 32 = 2 * 2 * 2 * 2 * 2

Work Shown

32 is divisible by 2: 32 = 16 * 2.
16 is divisible by 2: 16 = 8 * 2.
8 is divisible by 2: 8 = 4 * 2.
4 is divisible by 2: 4 = 2 * 2.
2 is not divisible by anything.

This calculation used this Prime Factorization Algorithm.


Solution by Factor any number
37, 53, and 157 are prime numbers.
108
Solved by pluggable solver: Factor any number
108 is NOT a prime number: 108 = 2 * 2 * 3 * 3 * 3

Work Shown

108 is divisible by 2: 108 = 54 * 2.
54 is divisible by 2: 54 = 27 * 2.
27 is divisible by 3: 27 = 9 * 3.
9 is divisible by 3: 9 = 3 * 3.
3 is not divisible by anything.

This calculation used this Prime Factorization Algorithm.


Solution by Factor any number
Note: If a number is prime, the only factors are 1 and the number.